# Spacetime Measurements with the Photon Ring

### v2

@@ Abstract
We explore the universal symmetries of the black hole photon ring in a wide range of non-Kerr spacetimes, including the Kerr-Newman, Kerr-Sen, Kerr-Bardeen, and Kerr-Hayward metrics. The demagnification exponent ($\gamma$) controls the size and flux scaling of higher-order images, which appear in the photon ring, the time delay ($\tau$) determines the timing of their appearance, and the rotation parameter ($\delta$) relates their relative orientations on the image plane. Our investigation reveals distinct responses of these critical parameters to black hole spin, generalized charge, and observer inclination: $\gamma$ is predominantly influenced by charge and spin, $\tau$ is strongly affected by inclination, especially for near-extremal black holes, and $\delta$ is highly sensitive to spin. Notably, we find that the time delay provides an independent constraint on shadow size for polar observers, while the rotation parameter facilitates metric-independent spin measurements. Specifically, for Kerr black holes, the total variation in $\gamma$, $\tau$, and $\delta$ across all possible inclinations is $\lesssim 10\%$, $\lesssim 20\%$, and $\lesssim 60\%$, respectively. By contrast, the Kerr shadow size varies by only $\lesssim 8\%$.
@@ I Introduction
Black holes (BHs) provide a rare and powerful laboratory to probe both gravitational physics and astrophysics simultaneously. Horizon-scale observations in the electromagnetic spectrum have recently become possible with the Event Horizon Telescope (EHT). These high-resolution, millimeter-wavelength observations of the supermassive BHs in Messier 87 (M87*) and in the Milky Way (Sgr A*) produced the historic first images of BH “shadows”, and are transforming our understanding of their environments and of the curved spacetime near event horizons [1, 2, 3]. The EHT images consistently display a ring-shaped feature known as “emission ring”, whose diameter closely traces the “critical curve” of the BH shadow boundary [4, 5, 6, 7, 8, 9]. Due to limited angular resolution, the EHT has not yet been able to measure the spins of these BHs – a fundamental property with significant implications for BH formation, growth, accretion dynamics, galaxy-scale jet formation and even evolution on galactic scales. This is further complicated by the weak dependence of the size of the BH shadow on spin and inclination, varying by only $\sim 8\%$ for Kerr BHs [10, 11].
General relativity (GR) predicts that the strong gravitational lensing of extended emission sources around BHs produces a direct image as well as higher-order images, with the latter collectively referred to as the “photon ring” [12, 13, 14, 15, 16]. The photon ring has not yet been resolved in the EHT images. For optically-thin emission, the photon ring comprises an infinite sequence of subrings – formed by photons that complete at least half of a polar or latitudinal orbit around the BH photon shell before reaching our sky. These subrings display a self-similar structure, and are indexed by the number of polar turning points or half polar-orbits completed en route to the observer, converging exponentially to the critical curve [16, 15]. Because photons travel along null geodesics of spacetime, these subrings carry distinct imprints of the near-horizon BH geometry, making them a “sweet spot” for studying the gravitational physics of BHs: They probe the near-horizon region, the leading-order subring remains observable at beyond-Earth baselines despite high redshift, and, importantly, they are minimally affected by the complex astrophysical processes in the accretion disk and jet base.
@@
Non-Kerr BHs, both in GR or modified theories of gravity, feature “deviation parameters” or “generalized charges” in addition to mass and spin. The observable effects of these deviation parameters are expected to manifest in the strong-field regime near the event horizon. The EHT measurements of the shadow sizes of M87* and Sgr A* have been widely utilized to place constraints on potential deviations from the Kerr metric [29, 9, 11, 30, 9, 31, 32, 33, 34]. However, relying on a single observable often results in a degeneracy between the measurements of spin and deviation parameters from the Kerr solution. A Kerr BH shadow of a given size can be very well explained by a non-Kerr BH with different spin and non-zero charge [35, 36].
@@
The prospect of observing photon ring in the near future has sparked significant interest [14, 15, 16, 17]. While photon rings for Kerr BHs are relatively well-understood in terms of their image structure, polarization, and Fourier visibility [17, 37, 38, 39, 40, 16, 16, 15, 18, 19, 41, 42, 43, 44], their morphology in non-Kerr spacetimes is not fully understood yet (see, however, Refs. [45, 39, 40, 46, 47]). In this paper, we aim to advance our current understanding of photon rings by addressing some key questions: How does the structure of the photon ring vary across different BH spacetimes? Can detection of the photon ring lead to robust BH spin measurement? And, most importantly, can joint measurements of the shadow size and specific photon ring features provide stringent constraints on deviations from Kerr spacetime?
@@
Building on Ref. [16], in our previous paper [48], we have described the three critical parameters ($\gamma_{\rm p},\tau_{\rm p},\delta_{\rm p}$) that govern the demagnification, time delay, and rotation, respectively, of successive photon subrings on the image plane in a broad class of stationary and axisymmetric spacetimes. In Ref. [48], we employed the Johannsen-Psaltis (JP) metric [49] – a four-dimensional, geodesically-integrable, spinning metric – to determine the critical parameters in terms of two metric functions and the observer’s inclination angle. [^1] The JP metric [49] is described by four independent metric functions $f,F,N,B$ in general. However, we have shown in Sec. II of Ref. [50] that the metric function $f$ can be absorbed into the “Mino time” and that the metric function $B$ can be eliminated via a radial coordinate transformation. Thus, null geodesics in the JP metric and the critical lensing parameters are completely determined by $F$ and $N$. This metric is widely adopted for studying strong-field effects in a theory-agnostic framework, as it can represent a broad range of non-Kerr BHs in GR and modified theories. To systematically investigate variations in photon ring morphology across different spinning spacetimes, we analyze specific BH solutions, including Kerr-Newman [51], Kerr-Sen [52], Kerr-Bardeen [53], and Kerr-Hayward [53]. These non-Kerr spacetimes are controlled by an additional “charge” parameter $Q$, arising from their underlying physical fields, in addition to the total mass $M$ and spin $a$. All of these metrics reduce to the Kerr metric in the limit of vanishing charge. Through this analysis, we quantitatively assess the potential of photon ring measurements to constrain strong-field deviations from the Kerr geometry, with the following key outcomes:
@@ II Photon Ring Critical Parameters
The Johannsen-Psaltis (JP) metric in Boyer-Lindquist coordinates, $x^{\mu}=(t,r,\vartheta,\varphi)$, is expressed as [49, 50]
@@
Note, however, that shifting the Carter constant in this way does not change the photon orbit. Eqs. (4) and (5) reveal that knowledge of the metric functions $N$ and $F$ is sufficient to determine the orbits of arbitrary photons in a JP spacetime (we can set $B=1$ through a change of $r$, as discussed above).
@@
where the sign ±$\pm_{\vartheta}$ corresponds to sign of the polar angular momentum $k_{\vartheta}$ at the location of the observer. Thus, the conserved quantities $(\xi,\mathscr{I})$ directly determine the impact parameters of the photon. BH shadow boundary curve or the critical curve for parameterized metric (1) has been extensively studied in literature [54, 55, 56, 57].
Strong gravitational lensing close to a BH can compel photons to move on bound spherical orbits. The radial velocity, $|\dot{r}|=\mathscr{R}$, and acceleration, $|\ddot{r}|=\partial_{r}\mathscr{R}/2$, of photons on such spherical null geodesics (SNGs) both must naturally vanish. This is sufficient to infer the impact parameters $(\xi_{\mathrm{p}},\mathscr{I}_{\mathrm{p}})$ of an SNG on a sphere of radius $r=r_{\mathrm{p}}$ as being
@@
where, in the above, $H(x)$ is the Heaviside step function. These critical parameters not only govern the nearly-bound photon orbits around the BH but also determine the scaling of higher-order subrings on the image screen. We have also introduced the hatted functions $\hat{G}$ which are given in terms of complete elliptic functions as follows [58, 16]
@@
with $K(k)$ the complete elliptic function of the first kind, $E^{\prime}(k)$ the derivative of the complete elliptic function of the second kind w.r.t. the elliptic modulus $k$, i.e., $E^{\prime}(k):=[E(k)-K(k)]/(2k)$, and $\Pi(n|k)$ the complete elliptic integral of the third kind, with $n$ denoting the elliptic characteristic. Our convention for the elliptic functions can be found in Ref. [48].
The prefactor of the term involving $\hat{G}_{\vartheta}$ in $\gamma_{\mathrm{p}}$ is the (Mino time) phase space Lyapunov exponent, $\kappa_{\mathrm{p}}=(\partial_{r}^{2}\mathscr{R}_{\mathrm{p}}/2)^{1/2}$. This Lyapunov exponent measures the radial instability of the SNGs, and governs the radial deviation, $\delta r(\lambda_{\mathrm{m}})$, of nearly-bound photon orbits from their bound counterparts (i.e., those having the same impact parameters),
@@
In the above, $\delta r(0)$ represents the initial radial coordinate distance between the two orbits. For photons on radially-perturbed orbits that are moving away from (towards) the SNG, one picks the positive (negative) sign in the equations above.
@@
Furthermore, the nearly-bound null geodesics that do appear on the image plane appear at different image plane polar angles, $\psi$, where
@@
Below, however, we elect to not use the wrapped rotation parameter since when averaged, $\langle\delta_{2\pi;\mathrm{p}}\rangle_{\psi}:=(1/2\pi)\int_{0}^{2\pi}\delta_{2\pi;\mathrm{p}}(\psi)\mathrm{d}\psi$, it introduces unnecessary wrapping artifacts (see right column in Fig. 7). Instead, $\langle\delta\rangle_{\psi}$ defined in Eq. (18) above will describe the mean rotation parameter. The appendix A discusses this issue in further detail. We also address there the impact on our conclusions of our choice to use the mean as a characteristic measure of the critical parameters on the image plane versus, e.g., the median. While the mean and median measurements of $\delta_{\rm p}$ agree at small inclinations, they differ significantly at large inclinations, where, unlike the mean $\delta_{\rm p}$, the median $\delta_{\rm p}$ remains nearly independent of inclination. This behavior can be partially attributed to the following reasoning: The portion of the photon shell contributing to the photon ring formation grows smoothly with the observer’s inclination – ranging from a single SNG, $r_{\rm p}^{0}$, for a polar observer to the entire photon shell, $r_{\mathrm{p}}^{-}\leq r<r_{\mathrm{p}}^{+}$, for an equatorial observer. Thus, at large inclinations, different segments of the photon ring, constructed from distinct SNGs with varying BL radii, experience differential demagnification, delay, and rotation. While the mean, as a statistical average, is sensitive to outliers (extremely high or low values) caused by prograde and retrograde SNGs near the equatorial plane, the median is not. Therefore, the mean $\delta_{\rm p}$ provides a more reliable statistical average than the median.
@@ III Variation of Critical Parameters with Observer Inclination and Non-Kerr Spacetimes
Figure 3: Variation of the fractional deviation in mean critical parameters – demagnification exponent (left column), time delay (center column), and rotation parameter (right column) – for Kerr-Newman BHs with charge and spin, as observed by an observer at a low inclination of $17^{\circ}$ (top row) and a high inclination of $89^{\circ}$(bottom row).
@@
Future observations, capable of resolving photon subrings, would be able to sample the critical parameters at multiple screen angles. For instance, resolved images of flares caused by orbiting hot-spots might enable measurements of all three critical parameters by correlating intensity fluctuations across successive ordered images [41]. In practice, observables would likely be derived from averages of events across multiple occurrences of various phenomena.
@@
Roughly motivated by these expectations, we study the variations in the $\psi-$averaged values of the critical parameters from their Schwarzschild values with varying spacetime geometries and observer inclinations. We discuss the impact of using the $\psi-$median values of the critical parameters as an alternative characteristic measure in Appendix B. The specific quantities we study are defined in Eq. (18) and Eq. (21). We represent our results through four main studies depicted in Figures 1, 2, 3 and 4 below.
@@
Figure 1 shows the variation in the mean values of critical parameters across different spacetimes as a function of observer inclination. Specifically, we present the fractional deviation of these mean values from their corresponding values in a Schwarzschild BH spacetime, emphasizing the influence of nonzero spin and charge with varying inclination. In this figure, the spin and charge of these BHs are fixed at intermediate values of $a=0.5M$ and $Q=0.4M$, respectively. The variation in all three critical parameters across BHs spacetimes – containing a variety of matter fields – is modest ($\lesssim 5\%$) for all observer inclinations. For the chosen spin and charge, across all studied non-Kerr spacetimes, the maximum deviations observed in the demagnification exponent, the time delay, and the rotation parameter from their corresponding Schwarzschild values are $\lesssim 10\%$, $\lesssim 5\%$, and $\lesssim 25\%$, respectively. This suggests that the demagnification exponent and the rotation parameter are more sensitive probes of the spacetime geometry than the time delay, independently of the observer’s inclination. Moreover, for a fixed spin and charge, different spacetimes exhibit distinct deviations in these critical values.
@@
To better understand which aspects of the spacetime each critical parameter is sensitive to, we allow the BH spin to vary in Fig. 2. We present our analysis for three representative spacetimes: Kerr BHs, Kerr-Newman BHs, and regular Kerr-Bardeen BHs. The BH charge is kept fixed to the previous value of $Q=0.4M$. Thus, in this figure, the vertical red lines correspond to the BH models shown in Fig. 1. Across the three spacetimes, the mean time delay, $\langle\bar{\tau}\rangle_{\psi}$, shown in the middle column, features complex behavior with varying BH spin and the observer inclination. Notably, across all three spacetimes and over the entire parameter space of ($a,\mathscr{i}$), we find $\langle\bar{\gamma}\rangle_{\psi}\leq 0$ and $\langle\bar{\delta}\rangle_{\psi}\geq 0$, while $\langle\bar{\tau}\rangle_{\psi}\leq 0$, except for high inclinations or near-extremal BHs. This indicates that a non-Schwarzschild BH can exhibit the same time delay as a Schwarzschild BH, yet differences in the demagnification exponent and rotation parameter can break this degeneracy, allowing these spacetimes to be distinguished. Nevertheless, the time delay, $\langle\bar{\tau}\rangle_{\psi}$, remains within $\approx 5\%$ of its Schwarzschild value, except in regions of large spins and high inclinations, as seen in the upper-right quadrants of the parameter space. The sharp increase in time delay at large spins and high inclinations – particularly for near-extremal BHs – arises from contributions from photon trajectories approaching the BH event horizon, where gravitational time dilation becomes significant. For an extremal ($a=M$) Kerr BH in particular, Eq. (13) can be used to show that photons with trajectories close to the horizon also remain close to the equatorial plane, and are, thus, only seen by observers located close to the equatorial plane. In contrast, for typical (non-extremal) BHs, the time delay is relatively insensitive to the observer’s inclination. Moreover, for small inclinations, the time delay varies by approximately 8% with spin – similar to the variation in shadow size (see Eq. (45) and Fig. 1 in Ref. [48]).
@@
We interpret the sensitivity of the rotation parameter to spin alone as a direct manifestation of its role in measuring the magnitude of frame-dragging, or gravitomagnetism, experienced by nearly-bound photon orbits. Notably, the first term in $\delta_{\rm p}$ in Eq. (14) varies linearly with spin. Furthermore, if such measurement yields a nonzero value in particular, $\langle\bar{\delta}\rangle_{\psi}\neq 0$ it would correspond to a measurement of frame-dragging close to a BH horizon, directly analogous to a measurement of the gravitomagnetic precession of the Gravity Probe B gyroscopes in the Earth’s gravitational field or that measured in the double pulsar system [62].
@@
This figure clearly demonstrates that the isocontours of $\langle\bar{\rho}\rangle_{\psi}$ and $\langle\bar{\tau}\rangle_{\psi}$ are nearly indistinguishable, effectively overlapping. Consequently, the mean time delay serves as an independent estimator of shadow size, particularly at low observer inclinations (see Eq. (45) and Fig. 1 of Ref. [48] for details on the time delay’s dependence on shadow size at low inclinations). In observations of M87*, where the jet orientation suggests a modest inclination angle relative to the observer, this result is particularly advantageous: a high temporal resolution measurement of time delays from transient events, such as flaring, offers an independent and complementary estimate of the shadow size. Furthermore, this figure illustrates the power of joint measurements of the photon ring critical parameters to pin down physical properties of astrophysical BHs, viz., their spin and charge. Joint measurements of either shadow size or time delay, along with the demagnification exponent and rotation parameter, give precise estimations of both BH spin and generalized charge.
Finally, to assess whether the time delay consistently provides a reliable measure of the shadow size, in Fig. 5, we replicate Fig. 4 but change the observer inclination to $\mathscr{i}=89^{\circ}$. It is immediately evident that both quantities, $\langle\bar{\rho}\rangle_{\psi}$ and $\langle\bar{\tau}\rangle_{\psi}$, exhibit distinct behaviors with varying BH spin and charge, and that they conspire to produce identical variations only for observers present close to the BH spin axis. We have discussed this latter feature in Ref. [48] (see Eq. (44) there). Note that two isocontours for $5\%$ deviation in time delay (solid curves) correspond to $\langle\bar{\tau}\rangle_{\psi}=\pm 0.05$.

### v3

@@ Abstract
We explore the universal symmetries of the black hole photon ring in a wide range of non-Kerr spacetimes, including the Kerr-Newman, Kerr-Sen, Kerr-Bardeen, and Kerr-Hayward metrics. The demagnification exponent ($\gamma$) controls the size and flux scaling of higher-order images, which appear in the photon ring, the time delay ($\tau$) determines the timing of their appearance, and the rotation parameter ($\delta$) relates their relative orientations on the image plane. Our investigation reveals that these critical parameters respond distinctly to variations in black hole spin, generalized charge, and observer inclination, establishing them as complementary probes of spacetime geometry: $\gamma$ is predominantly influenced by charge and spin, $\tau$ is strongly affected by inclination, especially for near-extremal black holes, and $\delta$ is highly sensitive to spin. Notably, we find that the time delay provides an independent constraint on shadow size for polar observers, while the rotation parameter facilitates metric-independent spin measurements. Specifically, for Kerr black holes, the total variation in $\gamma$, $\tau$, and $\delta$ across all possible inclinations and spins is $\lesssim 20\%$, $\lesssim 10\%$, and $\lesssim 60\%$, respectively. By contrast, the Kerr shadow radius varies by only $\lesssim 8\%$. A future joint measurement of these critical parameters—along with the black hole shadow size—will enable precise spacetime characterization, including measurements of the spin, inclination, and generalized charge.
@@ I Introduction
Black holes (BHs) provide a rare and powerful laboratory to probe both gravitational physics and astrophysics simultaneously. Horizon-scale observations in the electromagnetic spectrum have recently become possible with the Event Horizon Telescope (EHT). These high-resolution, millimeter-wavelength observations of the supermassive BHs in Messier 87 (M87*) and in the Milky Way (Sgr A*) produced the historic first images of BH “shadows”, and are transforming our understanding of their environments and of the curved spacetime near event horizons [1, 2, 3]. The EHT images consistently display a ring-shaped feature known as “emission ring”, whose diameter closely traces the “critical curve” of the BH shadow boundary [4, 5, 6, 7, 8, 9]. Due to limited angular resolution, the EHT has not yet been able to measure the spins of these BHs—a fundamental property with significant implications for BH formation, growth, accretion dynamics, galaxy-scale jet formation and even evolution on galactic scales. This is further complicated by the weak dependence of the size of the BH shadow on spin and inclination, varying by only $\sim 8\%$ for Kerr BHs [10, 11].
General relativity (GR) predicts that the strong gravitational lensing of extended emission sources around BHs produces a direct image as well as higher-order images, with the latter collectively referred to as the “photon ring” [12, 13, 14, 15, 16]. The photon ring has not yet been resolved in the EHT images. For optically thin emission, the photon ring comprises an infinite sequence of subrings—formed by photons that complete at least half of a polar or latitudinal orbit around the BH photon shell before reaching our sky. These subrings display a self-similar structure, and are indexed by the number of polar turning points or half polar-orbits completed en route to the observer, converging exponentially to the critical curve [16, 15]. Because photons travel along null geodesics of spacetime, these subrings carry distinct imprints of the near-horizon BH geometry, making them a “sweet spot” for studying the gravitational physics of BHs: They probe the near-horizon region, the leading-order subring remains observable at beyond-Earth baselines despite high redshift, and, importantly, they are minimally affected by the complex astrophysical processes in the accretion disk and jet base.
@@
Non-Kerr BHs, both in GR or modified theories of gravity, feature “deviation parameters” or “generalized charges” in addition to mass and spin. The observable effects of these deviation parameters are expected to manifest in the strong-field regime near the event horizon. The EHT measurements of the shadow sizes of M87* and Sgr A* have been widely utilized to place constraints on potential deviations from the Kerr metric [29, 9, 11, 30, 9, 31, 32, 33, 34]. However, relying on a single observable often results in a degeneracy between the measurements of spin and deviation parameters from the Kerr solution. A Kerr BH shadow of a given size can be very well explained by a non-Kerr BH with different spin and nonzero charge [35, 36].
@@
The prospect of observing photon ring in the near future has sparked significant interest [14, 15, 16, 17]. While photon rings for Kerr BHs are relatively well-understood in terms of their image structure, polarization, and Fourier visibility [17, 37, 38, 39, 40, 16, 16, 15, 18, 19, 41, 42, 43, 44], their morphology in non-Kerr spacetimes is not fully understood yet (see, however, Refs. [45, 39, 40, 46, 47, 48, 49]). In this paper, we aim to advance our current understanding of photon rings by addressing some key questions: How does the structure of the photon ring vary across different BH spacetimes? Can detections of the photon ring lead to robust BH spin measurement? And, most importantly, can joint measurements of the shadow size and specific photon ring features provide stringent constraints on deviations from Kerr spacetime?
@@
Building on Ref. [16], in our previous paper [50], we have described the three critical parameters ($\gamma_{\rm p},\tau_{\rm p},\delta_{\rm p}$) that govern the demagnification, time delay, and rotation, respectively, of successive photon subrings on the image plane in a broad class of stationary and axisymmetric spacetimes. In Ref. [50], we employed the Johannsen metric [51]—a four-dimensional, geodesically integrable, spinning metric—to determine the critical parameters in terms of two metric functions and the observer’s inclination angle.  [^1] The Johannsen metric [51] is described by four independent metric functions $f,F,N,B$ in general. However, we have shown in Sec. II of Ref. [52] that the metric function $f$ can be absorbed into the “Mino time” and that the metric function $B$ can be eliminated via a radial coordinate transformation. Thus, null geodesics in the Johannsen metric and the critical lensing parameters are completely determined by $F$ and $N$. This metric is widely adopted for studying strong-field effects in a theory-agnostic framework, as it can represent a broad range of non-Kerr BHs in GR and modified theories. This paper explores the astrophysical implications of these critical parameters. We first systematically analyze their variation with the spin and inclination of Kerr BHs and then quantitatively illustrate how deviations from the Kerr solution affect these parameters. To systematically investigate variations in photon ring morphology across different spinning spacetimes, we analyze specific BH solutions, including Kerr-Newman [53], Kerr-Sen [54], Kerr-Bardeen [55], and Kerr-Hayward [55]. These non-Kerr spacetimes are controlled by an additional “charge” parameter $Q$, arising from their underlying physical fields, in addition to the total mass $M$ and spin $a$. All of these metrics reduce to the Kerr metric in the limit of vanishing charge. Through this analysis, we quantitatively assess the potential of photon ring measurements to constrain strong-field deviations from the Kerr geometry, with the following key outcomes:
@@ II Photon Ring Critical Parameters
The Johannsen metric in Boyer-Lindquist coordinates, $x^{\mu}=(t,r,\vartheta,\varphi)$, is expressed as [51, 52]
@@
Note, however, that shifting the Carter constant in this way does not change the photon orbit. Equations (4) and (5) reveal that knowledge of the metric functions $N$ and $F$ is sufficient to determine the orbits of arbitrary photons in a Johannsen spacetime (we can set $B=1$ through a change of $r$, as discussed above).
@@
where the sign ±$\pm_{\vartheta}$ corresponds to sign of the polar angular momentum $k_{\vartheta}$ at the location of the observer. Thus, the conserved quantities $(\xi,\mathscr{I})$ directly determine the impact parameters of the photon. BH shadow boundary curve or the critical curve for parameterized metric (1) has been extensively studied in literature [56, 57, 58, 59]. Furthermore, the chaotic behavior of null geodesics and the resulting BH shadows are explored in the context of nonintegrable spacetimes, revealing complex and intriguing gravitational dynamics [60, 61, 62, 63, 64].
Strong gravitational lensing close to a BH can compel photons to move on a fixed BL radius orbit—bound spherical orbits. The radial velocity, $|\dot{r}|=\mathscr{R}$, and acceleration, $|\ddot{r}|=\partial_{r}\mathscr{R}/2$, of photons on such spherical null geodesics (SNGs) both must naturally vanish. This is sufficient to infer the impact parameters $(\xi_{\mathrm{p}},\mathscr{I}_{\mathrm{p}})$ of an SNG on a sphere of radius $r=r_{\mathrm{p}}$ as being
@@
where, in the above, $H(x)$ is the Heaviside step function. These critical parameters not only govern the nearly bound photon orbits around the BH but also determine the scaling of higher-order subrings on the image screen. We have also introduced the hatted functions $\hat{G}$ which are given in terms of complete elliptic functions as follows [65, 16]
@@
with $K(k)$ the complete elliptic function of the first kind, $E^{\prime}(k)$ the derivative of the complete elliptic function of the second kind with respect to the elliptic modulus $k$, i.e., $E^{\prime}(k):=[E(k)-K(k)]/(2k)$, and $\Pi(n|k)$ the complete elliptic integral of the third kind, with $n$ denoting the elliptic characteristic. Our convention for the elliptic functions can be found in Ref. [50].
The prefactor of the term involving $\hat{G}_{\vartheta}$ in $\gamma_{\mathrm{p}}$ is the (Mino time) phase space Lyapunov exponent, $\kappa_{\mathrm{p}}=(\partial_{r}^{2}\mathscr{R}_{\mathrm{p}}/2)^{1/2}$. This Lyapunov exponent measures the radial instability of the SNGs, and governs the radial deviation, $\delta r(\lambda_{\mathrm{m}})$, of nearly bound photon orbits from their bound counterparts (i.e., those having the same impact parameters),
@@
In the above, $\delta r(0)$ represents the initial radial coordinate distance between the two orbits. For photons on radially perturbed orbits that are moving away from (toward) the SNG, one picks the positive (negative) sign in the equations above.
@@
Furthermore, the nearly bound null geodesics that do appear on the image plane appear at different image plane polar angles, $\psi$, where
@@
Below, however, we elect to not use the wrapped rotation parameter since when averaged, $\langle\delta_{2\pi;\mathrm{p}}\rangle_{\psi}:=(1/2\pi)\int_{0}^{2\pi}\delta_{2\pi;\mathrm{p}}(\psi)\mathrm{d}\psi$, it introduces unnecessary wrapping artifacts (see right column in Fig. 7). Instead, $\langle\delta\rangle_{\psi}$ defined in Eq. (18) above will describe the mean rotation parameter. The appendix A discusses this issue in further detail. We also address there the impact on our conclusions of our choice to use the mean as a characteristic measure of the critical parameters on the image plane versus, e.g., the median. While the mean and median measurements of $\delta_{\rm p}$ agree at small inclinations, they differ significantly at large inclinations, where, unlike the mean $\delta_{\rm p}$, the median $\delta_{\rm p}$ remains nearly independent of inclination. This behavior can be partially attributed to the following reasoning: The portion of the photon shell contributing to the photon ring formation grows smoothly with the observer’s inclination—ranging from a single SNG, $r_{\rm p}^{0}$, for a polar observer to the entire photon shell, $r_{\mathrm{p}}^{-}\leq r<r_{\mathrm{p}}^{+}$, for an equatorial observer. Thus, at large inclinations, different segments of the photon ring, constructed from distinct SNGs with varying BL radii, experience differential demagnification, delay, and rotation. While the mean, as a statistical average, is sensitive to outliers (extremely high or low values) caused by prograde and retrograde SNGs near the equatorial plane, the median is not. Therefore, the mean $\delta_{\rm p}$ provides a more reliable statistical average than the median.
@@ III Variation of Critical Parameters with Observer Inclination and Non-Kerr Spacetimes
Figure 3: Variation of the fractional deviation in mean critical parameters—demagnification exponent (left column), time delay (center column), and rotation parameter (right column)—for Kerr-Newman BHs with charge and spin, as observed by an observer at a low inclination of $17^{\circ}$ (top row) and a high inclination of $89^{\circ}$(bottom row).
@@
Future observations, capable of resolving photon subrings, would be able to sample the critical parameters at multiple screen angles. For instance, resolved images of flares caused by orbiting hotspots might enable measurements of all three critical parameters by correlating intensity fluctuations across successive ordered images [41]. In practice, observables would likely be derived from averages of events across multiple occurrences of various phenomena.
@@
Roughly motivated by these expectations, we study the variations in the $\psi-$averaged values of the critical parameters from their Schwarzschild values with varying spacetime geometries and observer inclinations. We discuss the impact of using the $\psi-$median values of the critical parameters as an alternative characteristic measure in Appendix B. The specific quantities we study are defined in Eq. (18) and Eq. (21). We represent our results through four main studies depicted in Figs. 1, 2, 3 and 4 below.
@@
Figure 1 shows the variation in the mean values of critical parameters across different spacetimes as a function of observer inclination. Specifically, we present the fractional deviation of these mean values from their corresponding values in a Schwarzschild BH spacetime, emphasizing the influence of nonzero spin and charge with varying inclination. In this figure, the spin and charge of these BHs are fixed at intermediate values of $a=0.5M$ and $Q=0.4M$, respectively. The variation in all three critical parameters across BHs spacetimes—containing a variety of matter fields—is modest ($\lesssim 5\%$) for all observer inclinations. For the chosen spin and charge, across all studied non-Kerr spacetimes, the maximum deviations observed in the demagnification exponent, the time delay, and the rotation parameter from their corresponding Schwarzschild values are $\lesssim 10\%$, $\lesssim 5\%$, and $\lesssim 25\%$, respectively. This suggests that the demagnification exponent and the rotation parameter are more sensitive probes of the spacetime geometry than the time delay, independently of the observer’s inclination. Moreover, for a fixed spin and charge, different spacetimes exhibit distinct deviations in these critical values.
@@
To better understand which aspects of the spacetime each critical parameter is sensitive to, we allow the BH spin to vary in Fig. 2. We present our analysis for three representative spacetimes: Kerr BHs, Kerr-Newman BHs, and regular Kerr-Bardeen BHs. The BH charge is kept fixed to the previous value of $Q=0.4M$. Thus, in this figure, the vertical red lines correspond to the BH models shown in Fig. 1. Across the three spacetimes, the mean time delay, $\langle\bar{\tau}\rangle_{\psi}$, shown in the middle column, features complex behavior with varying BH spin and the observer inclination. Notably, across all three spacetimes and over the entire parameter space of ($a,\mathscr{i}$), we find $\langle\bar{\gamma}\rangle_{\psi}\leq 0$ and $\langle\bar{\delta}\rangle_{\psi}\geq 0$, while $\langle\bar{\tau}\rangle_{\psi}\leq 0$, except for high inclinations or near-extremal BHs. This indicates that a non-Schwarzschild BH can exhibit the same time delay as a Schwarzschild BH, yet differences in the demagnification exponent and rotation parameter can break this degeneracy, allowing these spacetimes to be distinguished. Nevertheless, the time delay, $\langle\bar{\tau}\rangle_{\psi}$, remains within $\approx 5\%$ of its Schwarzschild value, except in regions of large spins and high inclinations, as seen in the upper-right quadrants of the parameter space. The sharp increase in time delay at large spins and high inclinations—particularly for near-extremal BHs—arises from contributions from photon trajectories approaching the BH event horizon, where gravitational time dilation becomes significant. For an extremal ($a=M$) Kerr BH in particular, Eq. (13) can be used to show that photons with trajectories close to the horizon also remain close to the equatorial plane, and are, thus, only seen by observers located close to the equatorial plane. In contrast, for typical (nonextremal) BHs, the time delay is relatively insensitive to the observer’s inclination. Moreover, for small inclinations, the time delay varies by approximately 8% with spin—similar to the variation in shadow size (see Eq. (45) and Fig. 1 in Ref. [50]).
@@
We interpret the sensitivity of the rotation parameter to spin alone as a direct manifestation of its role in measuring the magnitude of frame-dragging, or gravitomagnetism, experienced by nearly bound photon orbits. Notably, the first term in $\delta_{\rm p}$ in Eq. (14) varies linearly with spin. Furthermore, if such measurement yields a nonzero value in particular, $\langle\bar{\delta}\rangle_{\psi}\neq 0$ it would correspond to a measurement of frame-dragging close to a BH horizon, directly analogous to a measurement of the gravitomagnetic precession of the Gravity Probe B gyroscopes in the Earth’s gravitational field or that measured in the double pulsar system [69].
@@
This figure clearly demonstrates that the isocontours of $\langle\bar{\rho}\rangle_{\psi}$ and $\langle\bar{\tau}\rangle_{\psi}$ are nearly indistinguishable, effectively overlapping. Consequently, the mean time delay serves as an independent estimator of shadow size, particularly at low observer inclinations (see Eq. (43) and Fig. 1 of Ref. [50] for details). In observations of M87*, where the jet orientation suggests a modest inclination angle relative to the observer, this result is particularly advantageous: a high temporal resolution measurement of time delays from transient events, such as flaring, offers an independent and complementary estimate of the shadow size. Furthermore, this figure illustrates the power of joint measurements of the photon ring critical parameters to pin down physical properties of astrophysical BHs, viz., their spin and charge. Joint measurements of either shadow size or time delay, along with the demagnification exponent and rotation parameter, give precise estimations of both BH spin and generalized charge.
Finally, to assess whether the time delay consistently provides a reliable measure of the shadow size, in Fig. 5, we replicate Fig. 4 but change the observer inclination to $\mathscr{i}=89^{\circ}$. It is immediately evident that both quantities, $\langle\bar{\rho}\rangle_{\psi}$ and $\langle\bar{\tau}\rangle_{\psi}$, exhibit distinct behaviors with varying BH spin and charge, and that they conspire to produce identical variations only for observers present close to the BH spin axis. Note that two isocontours for $5\%$ deviation in time delay (solid curves) correspond to $\langle\bar{\tau}\rangle_{\psi}=\pm 0.05$.